### v2

# AI’s Social Forcefield: Reshaping Distributed Cognition in Human–AI Teams

Abstract.
AI is not only a neutral tool in team settings; it actively reshapes the social and cognitive fabric of collaboration. We advance a unified framework of alignment in distributed cognition in human–AI teams—a process through which linguistic, cognitive, and social coordination emerge as human and AI agents co-construct a shared representational space. Across two studies, we show that exposure to AI-generated language shapes not only how people speak, but also how they think, what they attend to, and how they relate to each other. Together, these findings reveal how AI participation reorganizes the distributed cognitive architecture of teams: AI systems function as implicit “social forcefields.” Together, our findings highlight the double-edged impact of AI: the same mechanisms that enable efficient collaboration can also erode epistemic diversity and undermine natural alignment processes. We argue for rethinking AI in teams as a socially influential actor and call for new design paradigms that foreground transparency, controllability, and group-level dynamics to foster responsible, productive human-AI collaboration.
Human-AI teaming, collective attention, linguistics, entrainment, shared mental models, lexical alignment
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Some; 2025; Placeisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
AI is increasingly pervasive in modern workplaces (Brynjolfsson et al., 2025; Agarwal et al., 2023) and projected to have wide ranging impact on the future of work (BCG, 2024; AI Index, 2025). Yet many are increasingly concerned about its influence (Kennedy et al., 2025; Chayka, 2025). AI use is not isolated from other goings on as workers are embedded in teams and engaged in social interactions with human coworkers. The key thesis of this paper is that AI systems do not just passively assist human teams; they actively reshape the social and cognitive fabric of teamwork. Our work shifts the view of AI from a passive tool or assistant to an actor that actively influences cognitive alignment in human teams. We develop and test an unified framework of cognitive alignment in human–AI teams that links micro-level cognitive coupling (word reuse & shared phrasing) to higher-level collective cognition (shared focus of attention and shared understanding) and social cohesion & group identity. The framework specifies how AI-induced linguistic adaptation shapes shared attention and cognitive alignment, producing what we term AI’s social forcefield. Even when the AI is no longer present, linguistic traces continue to structure how humans coordinate with each other, shaping the persistent alignment of distributed cognition in human–AI teams.
This work is a detailed exploration how human-human communication is influenced by AI. It moves the conversation from “How can AI help humans?” to “How does AI rewire human collectives, and how do we design for that responsibly?” Our investigation considers cognitive alignment as an overarching, generative framework (Figure 1). We consider how lower-level influences cascade upward, via cumulative stabilization of alignment to shape distributed cognition in teams (Ocasio, 2011; Woolley et al., 2023; Riedl et al., 2021). Shared language reflects low-level cognitive coupling, enabling individuals to coordinate their representations in real time through mutual adaptation in speech (Clark, 1996; Sievers et al., 2024; Tausczik and Pennebaker, 2010). Linguistic entrainment in turn creates the representational scaffolding for shared attention and alignment of mental models (Clark, 1996; Kommol et al., 2023; Fusaroli et al., 2012; Hollan et al., 2000). That is, by naming things the same way, groups can sustain joint focus, and start seeing the task and each other through the same conceptual lens (Tomasello, 2009; Shteynberg, 2015; DeChurch and Mesmer-Magnus, 2010; Tausczik and Pennebaker, 2010). Shared mental models and attention in turn can amplify memories, goals, emotions, and behavioral strategies (Shteynberg, 2015; Clark and Brennan, 1991; Woolley et al., 2023), allowing groups to experience a sense of cognitive fluency and affective coherence (Hutchins, 1995); the feeling that “we understand each other” thus signals group identity and social cohesion (Tausczik and Pennebaker, 2010). The AI’s linguistic contributions penetrate upward through the layers of this theoretical framework—from phrasing, to shared focus of attention and shared mental models, to social cohesion and team identity—affecting the social and cognitive fabric of groups. Even when absent, the AI remains cognitively “in the loop.”
While recent work has made significant advances in understanding human-AI interaction (Rezwana and Maher, 2023; Riedl and Weidmann, 2025; Inkpen et al., 2023; Xiao et al., 2020; Bansal et al., 2019a; Branigan et al., 2011; Cowan et al., 2015; Wudarczyk et al., 2021; Zhang et al., 2023), these studies predominantly focus on dyadic relationships (one-user-one-AI) and immediate task performance (e.g., Zając et al., 2025; Vasconcelos et al., 2025; Arzberger et al., 2024; Ma et al., 2025). These studies often overlook the complexities of multi-member teams, and AI’s multifaceted effect on cognitive alignment across levels and beyond the direct interaction with AI. Prior work that is explicitly focused on teams, typically explores how teams use AI—or how AI can assist in teamwork—but often reported mixed results and high variability (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2023). We address these gaps by showing that AI creates a “social forcefield” that persists in human-human interaction, affecting cognitive alignment across multiple levels. Work often examines isolated AI effects (Wang et al., 2022) rather than interconnected cognitive alignment mechanisms in team settings (Cuevas et al., 2007). Even studies examining subtle AI influences, such as the placebo (Kosch et al., 2023) or ghostwriter effect (Draxler et al., 2024), investigate isolated effects during active AI engagement (Chaves et al., 2022), while we show how AI language spills over into subsequent human-human conversations and scaffolds cognitive alignment across multiple levels. As AI becomes an integral part of teamwork, addressing these challenges is critical to unlocking its potential to strengthen human-AI partnerships and guard against unintended consequences like homogenization. To address these gaps, we answer the following research questions:
  1. RQ1
    Does experimentally manipulating exposure to AI language causally alter team communication (language, attention, coordination) during collaboration, and does this influence spill over into subsequent human-only interaction?
  2. RQ2
    How robust are these effects across AI attributes (helpfulness and voice anthropomorphism), across interaction modalities (voice assistant and text-based chat), and across phases with and without active AI participation?
  3. RQ3
    Which social‑cognitive appraisals (e.g., trust, perceived intelligence, in‑group membership) predict the magnitude of these effects?
We provide evidence from two randomized controlled studies to show that exposure to AI-generated language shapes not only how people speak, but also how they think, what they attend to, and how they relate to each other. In Study 1, 20 teams of 3-4 individuals perform a 40-minute group task. Each team was randomly assigned a voice-only AI assistant that provided helpful or unhelpful information, and spoke in a human or robotic voice. Our analysis then explores (a) the AI’s effect on teams alignment on shared language, (b) teams shared focus of attention and mental models, and (c) group identity and cohesion. In Study 2, 97 individuals used ChatGPT 4o to answer customer service complaints. During this interaction, we exposed them to subtle wording differences by instructing the chatbot (via a randomly assigned system prompt) to suggest customer service responses that emphasized either warmth and empathy or formal adherence to company guidelines. Users then reflect on their approach to the task in a team-focused face-to-face debriefing interview. We then analyze overlap between subjects’ spoken words and the AI words they were exposed to by ChatGPT during the task. This design directly tests whether subtle variations in an AI assistant’s task framing and language leaves a linguistic and cognitive “fingerprint” that persists in human-human interaction beyond the AI’s direct involvement. Across both studies, a variety of data sources, and analysis approaches we find consistent evidence that AI shapes the cognitive alignment of human-human interaction, from the language people use when communicating with other humans, to their shared attention and mental models, and social cohesion and team identity.
In summary, we make the following contributions to our understanding of how AI influences cognitive alignment across multiple layers.
  1. (1)
    We show that AI significantly influences linguistic alignment, shared attention and mental models, and group identity and social cohesion.. This happens even though the AI assistant has limited control over team interactions (Study 1) or even when it is no longer present (Study 2), holds across different types of AI (voice assistant vs. text-based chat), quality levels (high, low, or naturalistic), or when subjects report not trusting it.
  2. (2)
    We reconceptualize AI as an implicit “social forcefield”: AI systems embedded in human teams inevitably shape human-human communication patterns, collective cognition, and shared identity—whether designers intend this or not. Unlike most studies that focus on one-on-one interactions (either human-to-human or human-to-AI), we examine how AI influences the communication and thinking of human groups beyond the direct interaction with the AI.
  3. (3)
    We highlight the double-edged impact of AI: It can help align language and distributed cognition, but also shapes collective cognition when not intended to—undermining natural alignment processes—and introduces homogeneity that reduces cognitive diversity.
  4. (4)
    We call for a new paradigm: AI will inevitably reshape human–human interaction. While its effects can be bounded through transparency and controllability, they cannot be fully predefined. We argue that effective human–AI teaming requires attention not only to functional performance, but also to the social and cognitive processes through which AI systems can enhance—rather than disrupt or homogenize—collective intelligence.
2. Background
Individual minds become coupled through communication and subtle linguistic cues can redirect what groups attend to and how they relate to each other. Our hypothesis is that AI assistants might not just be tools that provide information or automate tasks; they also affect the social and cognitive fabric of teamwork. Introducing AI in settings of human-human collaboration, might shift the way people talk (through linguistic entrainment) which in turn shifts what people pay attention and how they relate to each other. We conceptualize cognitive alignment as a generative framework that links micro-level linguistic coordination to macro-level collective outcomes (Figure 1). At the micro level, linguistic entrainment—individuals reusing words and phrasing introduced by the AI—creates a shared representational substrate. This alignment in language enables meso-level convergence in attention and shared mental models, as teams begin to focus on the same issues and structure the task through similar conceptual lenses. Over time, such cognitive coherence gives rise to macro-level social integration, expressed through in-group identification (e.g., inclusive pronouns like “we”). In this view, cognitive alignment is not a collection of parallel effects but a cascade of interdependent processes through which AI-introduced language shapes how people think, what they attend to, and how they relate to one another. This framework extends distributed cognition from being computer mediated and shaped by digital objects (Hollan et al., 2000) into the social domain of human–AI teams.
Figure 1. AI exerts a “social forcefield,” affecting the cognitive alignment in human groups through a multi-level model, from Micro-level coupling of linguistic entrainment, through meso-level alignment of attention and shared mental models, to macro-level alignment through relational and social integration.:
2.1. Shared Language
On the lowest level of our framework sits shared language. Shared language reflects the micro-level cognitive coupling that allows individuals to align their representations as they adapt to each other’s speech in real time. Lexical alignment occurs when conversational partners converge on similar words or phrases during interaction (Wachsmuth et al., 2013). Over time, these repetitions form “conceptual pacts,” which are agreements on how to refer to things within a group (Brennan and Clark, 1996). Alignment on shared language reduce cognitive effort; without shared language communication would be more difficult and increase processing demands (Metzing and Brennan, 2003). Shared language therefore helps groups coordinate attention, conserve mental resources, and collaborate more efficiently (Kalociński et al., 2018).
2.2. Shared Attention and Shared Mental Models
Shared attention reflects the meso-level convergence of focus, through which individuals align what aspects of the task are salient and meaningful, sustaining a common frame of reference for joint action. Language carries shared representations and structuring how people think and act together. Shared language alignment is thus the linguistic foundation of collective attention which shows priorities and intentions (Tausczik and Pennebaker, 2010). Human attention is the ability to actively process specific information while tuning out other, less relevant details (Bundesen, 1990). Attention is limited in both capacity and duration (Duncan, 2016), making its effective coordination essential for high performance (Duncan, 2013). In group contexts, individuals align their attention with others, including what they perceive, remember, and prioritize (Kampis and Southgate, 2020). Groups achieve high collective attention when members co-attend to the same object or event, creating a shared psychological experience that shapes motivation, judgment, and behavior (Shteynberg, 2015; Woolley et al., 2023). Collective attention thus enables teams to coordinate activities and sustain joint action (Mayo and Woolley, 2021).
Collective attention is often studied through multiple lenses: the temporal focus of groups on shared information (e.g., social media activity) (Weng et al., 2012), synchrony in behaviors like facial expression or vocal rhythm (Woolley et al., 2023), and, crucially, the alignment of language (Sievers et al., 2024). Language can play a central role because it links cognitive processes to shared meaning, aligning beliefs and neural activity across team members (Nagai et al., 2003; Kaplan and Hafner, 2006). By shaping how problems are described and discussed, collective attention also influences abstract intentions, plans, and strategies (Hawkins et al., 2023), which in turn can affect creativity and performance (Laureiro-Martinez et al., 2023; Bergenholtz et al., 2023).
Together with shared attention, shared mental models are the collective cognitive structures that teams develop to interpret and predict their environment (DeChurch and Mesmer-Magnus, 2010). They capture a group’s shared understanding of “what is happening, what is likely to happen next, and why it is happening” (Mohammed et al., 2010, p. 879). Such models improve motivation, coordination, and performance, particularly when tasks require mutual dependence on diverse skills and expertise (Lewis, 2003). Their development relies on the temporal alignment of attention and the establishment of shared language, which together provide the building blocks for joint understanding (Tollefsen et al., 2013; DeChurch and Mesmer-Magnus, 2010; Mohammed et al., 2010). That is, the shared lexicon on the micro-layer scaffolds shared understanding of the broader context of work through the development of shared mental models and joint focus of attention.
2.3. Group Identity and Social Cohesion
Social cohesion and team identity emerge from this cognitive coherence, as smooth coordination and shared understanding create a sense of mutual predictability and belonging—turning cognitive alignment into social alignment. Language can signals collective identity through the use of function words that require mutual understanding, such as pronouns (Swol et al., 2021; Sievers et al., 2024). Using pronouns only makes sense when speakers share an understanding of who the pronouns refer to. Groups with frequent pronoun use have a higher level of trust in their teammates and strong social cohesion (Tausczik and Pennebaker, 2010). The level of social cohesion and group identity is crucially dependent on who is considered part of the group. Teams often differentiate between in-group and out-group members, which alters how much alignment occurs and how shared mental models develop (Brennan and Hanna, 2009; Liberman et al., 2018). When an agent—such as an AI system—is treated as a teammate, it may facilitate greater alignment and integration of perspectives (Wang et al., 2021), but if it is seen as an outsider, alignment may be weaker or even resisted (Xie, 2023; Mahmud et al., 2023). Whether AI is integrated into the collaborative fabric of a team or remains at its margins is critical to understand its impact on the overall social fabric of the team (Groom et al., 2011; Nass and Moon, 2000).
In our framework, social cohesion and team identity represent macro-level manifestations of cognitive alignment. We measure them through surface-level linguistic signals—inclusive pronoun use—an approach that is theoretically grounded in the view that language both reflects and enacts shared cognitive and social structures (Hutchins, 1995; Clark, 1996). A substantial body of work demonstrates that function-word choices, particularly pronouns, provide reliable markers of interpersonal alignment, collective identity, and cohesion (Tausczik and Pennebaker, 2010; Gonzales et al., 2010). From a distributed cognition perspective, such linguistic traces reveal the degree to which a team members inhabit a shared representational space—a precondition for successful coordination and collective agency (Pickering and Garrod, 2014; Hutchins, 1995; Searle, 1995).
3. Related Work
Our research is most closely related to work on (a) human-AI teaming and (b) alignment on shared language between humans and artificial systems.
3.1. Human-AI Teaming
The growing importance of teamwork (Wuchty et al., 2007) has converged with rapid advances in artificial intelligence to fuel the development of collaborative human–AI teams (Malone and Bernstein, 2015; O’Neill et al., 2020). The vision is that such teams can leverage the complementary strengths of humans and AI, thereby transcending the limitations of human-only groups (Brynjolfsson et al., 2018; Vaccaro et al., 2024).
3.1.1. From Task Performance to Cognitive Alignment
Most research on human-AI interaction has emphasized performance outcomes, focusing on how AI systems automate processes, augment human decision-making, or generate novel solutions (e.g., Dell’Acqua et al., 2023; Chen and Chan, 2024; Meincke et al., 2024). Experimental setups often compare human-only performance to human-AI dyad performance, with measures such as speed, accuracy, and output quality. For example, studies in medical imaging show that AI can improve diagnostic accuracy when combined with radiologists’ judgments (e.g., Agarwal et al., 2023; Zając et al., 2025), and that AI collaboration can enhance decision-making performance (e.g., Fügener et al., 2022). However, results are inconsistent: some studies report performance gains, while others find disruption or reduced effectiveness (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2023). This variability highlights the need to look beyond outcomes and investigate the mechanisms that shape collaboration, such as collective attention and cognitive alignment.
Current AI systems are usually optimized for task success rather than for supporting group dynamics like shared focus of attention or collective cognition. This narrow focus can inadvertently disrupt multi-human collaboration, for example, by weakening a team’s ability to sustain collective attention. Research has made significant advances in shedding light on the important of alignment between humans understanding the AI and the AI understanding humans (Bansal et al., 2019a, b; Paleja, 2022; Alipour et al., 2021; Bansal et al., 2019a; Fügener et al., 2022). Much less is known about the indirect “spillover” from AI to human-human dynamics beyond the direct interaction with the AI.
3.1.2. Human-AI Alignment and Ethical Concerns
An emerging body of work explores how humans adapt linguistically and cognitively to AI. Spillner and Wenig (2021) showed that users exhibit lexical alignment with chatbots in digital companion settings, while Li and Song (2025) provided a framework for bidirectional alignment, emphasizing both human-to-AI and AI-to-human processes. However, this work generally focuses on dyadic interactions, leaving open the question of how AI-generated language influences dynamics in larger human groups and beyond the direct interaction with (or presence of) the AI system.
At the same time, researchers have raised ethical concerns about over-adaptation.  van der Rijt et al. (2025) caution that AI mimicking human language too closely can blur accountability and reinforce social biases. Relatedly, work on conversational agents in service encounters demonstrates how AI can subtly shape human communication practices (Van Pinxteren et al., 2020). These findings point to the broader social consequences of AI alignment, highlighting the importance of studying how AI reshapes collaboration beyond immediate task utility, and point to crucial ethical concerns of that arise when humans adapt to AI and/or AI is explicitly designed to steer such adaptation one way or another.
3.1.3. Toward Multi-Member Dynamics
Other studies have begun to explore how AI systems can enhance group dynamics more directly. For example, shared control between humans and robots allows teammates to combine complementary strengths, improving decision quality and coordination (Music and Hirche, 2017). Similarly, AI agents that scaffold meta-cognitive processes, such as mapping “who knows what” in a group, can strengthen shared mental models and collaboration (Westby and Riedl, 2023). These designs highlight the potential of AI to support not just task outcomes but also the cognitive and social alignment that underpins effective teamwork (Haindl et al., 2022; Bendell et al., 2021; Stowers et al., 2021; Ezer et al., 2019).
Yet despite these advances, most studies remain focused on human-AI dyads (e.g., Bansal et al., 2019a, b; Wang et al., 2019; Fügener et al., 2022; Alipour et al., 2021), leaving fundamental questions about how AI participation alters multi-member team dynamics (Dennis et al., 2023). Collective attention is not simply the sum of individual attention but emerges from multi-agent coordination among teammates and AI (Demir et al., 2019; Harrison et al., 2003). Prior research suggests these dynamics can be unpredictable: if one person adapts smoothly to the AI while another resists, overall team coordination may falter (Balachandar et al., 2019; Reichel et al., 2018).
Guingrich and Graziano (2024) hypothesized that AI-generated language could spill over into human–human collaboration even after the AI is removed, but provided no empirical evidence. Our work directly investigates this proposition, showing how linguistic entrainment to AI reshapes communication within human teams and leaves durable traces in group interaction.
3.2. Alignment on Shared Language in Human-AI Interaction
When people talk, they co-construct shared terminology to maintain common ground (Branigan et al., 2011; Garrod and Anderson, 1987). With AI, the dynamics differ: users often adapt to the system’s vocabulary to minimize errors, especially if they perceive the AI as less capable of adapting to them (Branigan et al., 2011; Hu et al., 2014; Weatherholtz et al., 2014). Evidence is mixed: some studies suggest stronger alignment with AI than with human partners (Branigan et al., 2011), while others find the opposite (Wudarczyk et al., 2021; Brandstetter et al., 2017). Importantly, alignment with AI sometimes persists beyond the interaction, influencing subsequent human linguistic choices (Brandstetter et al., 2017). Still, most of this work focuses on dyads, leaving unanswered how alignment unfolds in mixed human-AI groups beyond the direct interaction with AI and its broader consequences beyond just language alignment. Gingrich (Guingrich and Graziano, 2024) speculates about AI spillover into human–human teamwork but offers no empirical demonstration. Our study addresses this gap by showing how alignment with AI can leave lasting linguistic and cognitive traces in human group interactions.
3.3. Bridging to Our Research Questions
This review or related work highlights three primary gaps. First, research on human-AI collaboration has focused too narrowly on task outcomes, neglecting the cognitive and social processes that underpin teamwork. Second, existing studies of human-AI alignment are limited to dyadic interactions (one-human-one-AI) and situations of direct AI interaction, leaving open questions about how AI spills over to multi-member teams and situations in which the AI may not be present. Third, foundational work in human communication shows that alignment and entrainment are central to cognitive alignment, suggesting that AI could significantly reshape these processes. Past work has focused on isolated effects without exploring a broader framework of scaffolded effects across multiple layer of alignment. Building on these insights, our work investigates whether and how AI affects cognitive alignment across different layers in human-human interaction, whether alignment with AI language persists when AI is no longer present, and hos it cascades into higher-order cognitive processes including shared attention, mental models, and group identity and cohesion.
4. Study 1 - Team-Based Collaborative Problem-Solving
4.1. Study Design and Methods
The goal of our first study is to examine how an AI assistant might influence the cognitive alignment in human teams. For this purpose, we recruited 20 teams of three or four participants (69 subjects total) to work together on a collaborative problem-solving task via video conferencing. Our study is designed as a randomized controlled trial with a 2 $\times$ 2 between-subjects design. We manipulated the information provided by the AI (helpful vs. unhelpful) and voice (human-sounding vs. robotic-sounding) of the AI assistant. After completing the task, participants answered a survey about their perceptions of the task, the team, and the AI assistant (see Appendix A1 for details).
4.1.1. Experimental Task
Participants collaborated on the “Cursed Treasure Puzzle,” a pattern-recognition challenge that requires identifying hidden rules and combining them into a final solution. [^1] Task details available at original source at https://jaylorch.net/puzzles/CursedTreasure/ The puzzle is ideally suited to study collaborative problem-solving in human-AI settings: it requires teams to jointly infer hidden rules under uncertainty, coordinate distributed reasoning, manage iterative hypothesis and testing loops which are reflective of many real-world engineering and research tasks, and integrate partial insights across interdependent subproblems into a shared solution. Furthermore, the task requires communication economy to manage information overload and negotiation of shared understanding given its relies on unfamiliar symbols with no standard names (Clark and Wilkes-Gibbs, 1986; Reagans et al., 2023). These properties make it especially suited to studying emergent coordination, communication, and collective sensemaking—all core components of collaboration (additional justifications for these design choices are provided in Appendix A1). Teams had 40 minutes to work on the puzzle and received one AI-provided clue per milestone.
Figure 2. Three chests from the Cursed Treasure puzzle. A treasure chest is cursed with the “octopus curse” if the gems are separated but not cursed if the gems are touching.:
4.1.2. AI Assistant (Experimental Intervention)
The “AI Puzzle Master” was introduced as a voice-only assistant that provided pre-recorded audio messages at fixed intervals (see recordings here). “Puzzle Master” (sometimes “game master” or “MC”) is a widely used in contexts such as logic puzzle hunts, escape rooms, and Dungeons & Dragons–style games for the neutral facilitator who manages the task and provides feedback, and does not inherently imply social authority or evaluative power. In the helpful condition, messages provided relevant hints (e.g., “The octopus rule relates to the separation between the gems”). In the unhelpful condition, the messages gave misleading or irrelevant clues (e.g., it restarted obvious information about the task that was already known from the tasks instructions). The voice of the AI assistant was manipulated between human (recording of a professional female voice actor) and robotic-sounding (text-to-speech). An external panel validated that the robotic voice was perceived as non-human (see Appendix A1 for details). This controlled design allowed us to isolate the causal effects of AI messages on team processes, while keeping the AI’s role predictable and consistent across groups.
4.1.3. Measuring Linguistic Alignment
We analyzed team conversations to study how groups adopted (or resisted) the AI’s language. Using the full transcripts of all team communication (see Appendix A1 for details on recording and pre-processing steps) we developed a set of terms to track salient objects of the task (i.e., five core curses, plus peripheral terms such as “symbols” and “gems”). Each time a participant referred to a tracked object, we recorded whether they used the AI’s term or another term. Lexical alignment was calculated as the proportion of mentions using the AI’s term out of all mentions for that referent. Values range from 0 (never using the AI’s term) to 1 (always using the AI’s term). This measure allowed us to track how alignment with the AI’s language evolved during the puzzle (see Figure 3).
Figure 3. Illustration of lexical alignment on the AI assistant’s language in one example team.:
4.1.4. Survey Measures
After the puzzle, participants completed surveys on three topics:
  1. (1)
    Shared Mental Models: six-item scale on strategy agreement (Johnson et al., 2007; Mohammed et al., 2010).
  2. (2)
    Trust & AI Perceptions: six items on usefulness and reliability of the AI assistant (Pan and Steed, 2016).
  3. (3)
    AI as a Team Member: single item asking whether the AI was perceived as part of, or leading, the team.
4.1.5. Analysis Method
To estimate the causal effect of AI interventions on language use, we applied difference-in-difference regression models (Angrist and Pischke, 2009). The method compares changes in term usage for objects mentioned by the AI against control objects not mentioned by the AI in the same segment. The causal effect is identified by the inclusion of fixed effects for object, team, and time segment were included, with clustered standard errors at the team level (see full model details and robustness checks in Appendix A1). The analysis captures the causal effect of AI induced changes in language accounting for baseline likelihood of language changes, shifts in topic salience over time, and teams’ overall volume of communication.
4.1.6. Participants
We recruited 69 participants (59% female, mean age 21.7, 96% native or full English proficiency) in the United States. Teams of three or four were randomly assigned across conditions (Table 1). Each session lasted 90 minutes, and participants received \$20/hour in Amazon gift cards.
Treatment Teams Individuals Female Mean Age (SD) Native/Full Proficiency
Human/Helpful 5 17 65% 22.2 (4.56) 100%
Human/Unhelpful 5 17 47% 21.9 (3.37) 100%
Robotic/Helpful 5 18 56% 20.9 (2.78) 83%
Robotic/Unhelpful 5 17 71% 21.8 (3.62) 100%
All 20 69 59% 21.7 (3.67) 96%
Table 1. Participant demographics. English fluency was self-reported as Native, Full, or Proficient.
4.2. Results
This section reports the main findings of Study 1: how the AI assistant influenced the collective attention of teams, both in terms of what they focused on and how they talked about it. We present high-level results here and refer readers to Appendix A1 for detailed case studies, additional survey analyses, and evidence on underlying social and cognitive mechanisms.
4.2.1. Linguistic Alignment
The AI influenced how teams communicated about the objects of discussion, as reflected in the specific word choices they made. Teams generally adopted the AI’s terms after its interventions (Table 2, Model 1: $\beta=0.06$, $p<0.01$), but this effect varied by condition. Alignment was strongest in the Human-Helpful condition and weakest in the Robotic-Unhelpful condition (Table 2).
Dependent Variable: Entrainment on AI Language
Core Terms Peripheral Terms
(1) (2) (3) (4) (5)
After AI Intervention $0.06^{**}$ $0.14^{***}$ $0.10^{**}$ $0.15^{***}$ $0.16^{*}$
$(0.02)$ $(0.03)$ $(0.03)$ $(0.03)$ $(0.07)$
After $\times$ Robotic Voice $0.11^{*}$
$(0.04)$
After $\times$ Unhelpful 0.08-0.08 $0.27^{**}$
$(0.05)$ $(0.09)$
After $\times$ Human Unhelpful 0.04-0.04
$(0.06)$
After $\times$ Robotic Helpful $0.09^{**}$
$(0.03)$
After $\times$ Robotic Unhelpful $0.16^{*}$
$(0.06)$
Controls
   Referent Mentions $1^{\dagger} 0.00-0$$^{\dagger} 0.0$$1^{\dagger} 0.0$$1^{\dagger} 0.0$00.00
$(0.00)$ $(0.00)$ $(0.00)$ $(0.00)$ $(0.00)$
   Elapsed Time $0.00$ 0.00-0.00 $0.00$ 0.00-0.00 0.00-0.00
$(0.00)$ $(0.00)$ $(0.00)$ $(0.00)$ $(0.00)$
Num. obs. 1,995 1,995 1,995 1,995 713
Adj. R2 $0.70$ $0.70$ $0.70$ $0.70$ $0.79$
Fixed Effects: Team 2020 2020 2020 2020 2020
Fixed Effects: Object 55 55 55 55 22
Fixed Effects: Segment 77 77 77 77 77
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$; $^{\dagger}p<0.1$
Table 2. Difference-in-difference analysis of causal effect of AI intervention on teams’ alignment with AI language. The model accounts for repeat observations through inclusion of fixed effects and standard errors clustered on team-level. Note that the main effect for Robotic Voice and Unhelpful cannot be estimated as it is collinear with the team-level fixed effect.
Why were effects stronger for Human-Helpful? Our survey responses help explain this pattern. Participants rated helpful AIs as more trustworthy and intelligent ($p<0.001$), and they were more likely to see a human-sounding AI as part of their team. Language data further reinforced this: use of the pronoun “we” was highest in the Human-Helpful condition, suggesting the AI was integrated into collective identity and given higher status. In these cases, alignment seems to reflect not only informational influence but also a social process of treating the AI like a valued teammate.
Why the “flipped” effect for peripheral terms? An unexpected effect emerged for peripheral referents, which we define as puzzle elements that were part of the discussion but not essential for finding the solution (Table 2, Model 5). These elements include descriptive or secondary labels (e.g., “symbols,” “gems”) that teams could talk about without directly affecting their progress on the main milestones. In unhelpful conditions, teams resisted the AI’s terminology for core concepts but showed a higher likelihood of adopting them for peripheral ones. This likely reflects a low-stakes channel of influence: adopting the AI’s language for non-essential elements carried little cost and reduced cognitive effort, consistent with prior findings on linguistic alignment in low-stakes contexts (Vasconcelos et al., 2023). In contrast, for core concepts, excluding the AI’s language may have helped teams reinforce their own shared understanding and social cohesion.
4.2.2. Shared Attention & Alignment of Mental Models
The AI assistant had a consistent impact on what teams discussed, shaping their collective focus of attention. Teams were significantly more likely to talk about aspects of the task immediately after the AI introduced it (Figure 4). Using the same difference-in-difference analysis approach confirmed this effect ($\beta=0.35$; $p<0.001$), showing that the AI reliably seeded new discussion topics, regardless of whether its clues were helpful or unhelpful, or whether its voice was human or robotic.
Figure 4. AI affects the shared focus of attention. Teams talk more about the object mentioned by the AI in the segment immediately after its interjection than at other times. Counts include all mentions of tracked objects, regardless of which terms are used.:
The AI also had significant effect on teams shared mental models (one-way ANOVA $F(3,65)=5.5$; $p=0.041$). We find a consistent level of alignment in shared mental models in both helpful conditions, while the unhelpful conditions exhibit diverging patterns (Figure A2). Mental models are less well aligned (i.e., have higher standard deviation) in the Human-Unhelpful condition, while we find the highest level of alignment in the Robotic-Unhelpful condition. The combination of evidence from behavioral measures of group interaction (lexical alignment) and participants’ self-reflection of their group’s process (shared mental models) offers particularly compelling support for the AI assistant’s influence on cognitive alignment (DeChurch and Mesmer-Magnus, 2010; Lix et al., 2022).
4.2.3. Group Identity and Social Cohesion
Finally, we explore explore the AI’s effect on group identity and social cohesion through the use of first- and third-person pronouns (Figure 5; see Appendix A1 for additional method details). Teams in the Human-Helpful and Robotic-Helpful conditions showed a moderate level of pronoun use (a mean of 281 and 269 pronouns per team, respectively). Teams in the Robotic-Unhelpful treatment used considerably more pronouns (mean of 297 per team) indicating much stronger group identity social cohesion, while teams in the Human-Unhelpful treatment used significantly fewer (mean of 223 per team), suggesting little cohesion. In particular “we” is the most frequently used pronoun in the Human-Helpful condition but is used rarely in the Human-Unhelpful condition. This pattern suggests that the presence of an unhelpful, yet human-sounding AI assistant disrupted the natural cognitive alignment in teams and undermined their social cohesion. In the Robotic-Unhelpful condition, the unnatural voice of the AI makes it easier for the human participants to exclude it from their in-group thus strengthening their social cohesion. The failure to exclude the disruptive AI agent also explains the poor alignment of mental models in the Human-Unhelpful condition.
4.2.4. Collective Cognition as Multi-Level Alignment
Taken together, these results show that AI assistants do shape collective cognition not through a single mechanism, but through multi-level alignment. AI induces linguistic alignment of word choices, which further scaffolds shared attentional focus and alignment of mental models, and shifted perceptions of group membership. From a dynamical systems perspective, this fits with theories of human interaction as self-organizing synergies across multiple alignment channels, including joint attention, linguistic entrainment, social identity, and shared intentionality (Hultman et al., 2022; Gallotti et al., 2017). It is important to realize that even a scripted, non-interactive AI can perturb this alignment system, redirecting focus, reshaping language, and altering the social boundaries of teamwork.
Figure 5. Pronoun use indicates strong team cohesion in the Robotic-Unhelpful condition (based on successful exclusion of AI), yet disrupted team cohesion in the Human-Unhelpful condition (based on unsuccessful exclusion of AI). Bars show count of each pronoun across all teams in each treatment.
5. Study 2 - Customer Service Encounters
We complement our first study with a second preregistered [^2] https://osf.io/quj2d/?view_only=9299fa86594b487c9c356e1a32dbe6f7 and IRB approved study. This study uses a minimal design with a state-of-the-art interactive chatbot. It directly tests whether subtle variations in an AI assistant’s framing and language can leave a cognitive and linguistic “fingerprint” that persists when people later communicate with other humans—even when the AI is no longer present. The study design randomly manipulates the AI’s system prompt and measures how much subjects spontaneously reuse AI words in a human-led debriefing. This experimental manipulation of linguistic exposure thus isolates the causal pathway through which AI priming reshapes human language and thought.
Study Design and Task
To achieve high ecological validity, we designed a study around a realistic task—AI assisting customer support agents who perform text-based chat support to resolve customer complaints and technical problems. AI is already widely used for such tasks (BCG, 2024). One prominent field study in a Fortune 500 company showed that AI is highly effective supporting customer service agents, reporting an 14% average increase in productivity (Brynjolfsson et al., 2025). Our study had two phases. First, users interacted with a state-of-the-art chatbot (ChatGPT 4o) (Hurst et al., 2024) to write answers to five customer service complaints across different domains (including technical trouble shooting, online shopping assistance, and reservation booking; one example service complaint was “Every time I try to watch a movie, it buffers endlessly”). The AI was a fully functional ChatGPT 4o assistant. Subjects could interact with it like with any chatbot system, write any prompt they liked, for as many rounds as they wanted. Subjects interacted with the AI on a study-provided laptop computer. Users were randomly assigned to one of two AI conditions. Through the system prompt we instructed ChatGPT to either suggest customer service responses that emphasized warmth and empathy or formal adherence to company guidelines (see full system prompts in Appendix A2). The system prompt manipulation subtly changed which words subjects were exposed to by the AI. We term these exposure words and they provide the key experimental manipulation and basis for our analysis. We recorded all user-written prompts and ChatGPT responses (i.e., exposure words), as well as the ultimate response to the customer complaint through a custom user interface implemented via Qualtrics. Each conversation started with ChatGPT offering a suggested draft response to the customer complaint. Subjects wrote on average one additional prompt per customer complaint to refine the initial draft response generated by the chatbot ($SD=1$). The typical interaction was two prompts long. We find no significant different in engagement between the two treatment conditions (two sample t-test, $t=0.843$, $df=92.988$, $p=0.401$).
After the interaction with the AI, subjects participated in an in-person, face-to-face debriefing interview with an experimenter that lasted about five minutes (SD: 108 seconds). The debriefing interview questions were exclusively focused on the tasks—not the AI. To capture not only how the AI shapes how people talk, but also how they imagine acting in the future—their cognitive-behavioral approach to the task—questions were specifically designed to elicit subject’s process, goal, or approach to the task. For example, one question from the interview guide was “If you were working on a real customer service team, what kind of approach would you suggest the team agree on for answering complaints?” A common follow-up question explicitly elicited subject’s approach to the task and behavioral intention “What’s the first thing you usually focus on [when answering customer service complaints]?” In this way, our measure of overlap captures not only linguistic entrainment but also what aspects of the task are salient and meaningful to people, thus capturing important aspects of cognitive alignment through which groups sustain a common frame of reference for joint action (Shteynberg, 2015; Clark, 1996; Clark and Brennan, 1991; Tomasello, 2009; Hutchins, 1995).
Furthermore, interviews were explicitly framed as team-interaction questions. That is, the conversation was focus on collective group-level constructs, framed around how teams that the subject is part of: the collective unit of the customer service team (see quote above) (Mathieu et al., 2022). This form of “we” narratives are not only expressions shared language, mutual belief, and group identity (Gallagher and Tollefsen, 2019), they help construct and maintain social relations (Clark, 1996). Saying “we” instead of “I” positions the speaker as part of a collective agent and enacts a shared intentional frame. This framing was highly effective and 35% of utterances included the words “we” or “team” (example response from one subject: “[customers] should be let know that we are working on the problem and we resolve it as soon as possible and the team should be assigned as soon as possible to work on it, yeah”—emphasis ours). The conversation thus also reflects important social cohesion and team identity.
Experimenters were blind to the treatment condition and exposure words that subjects experienced during their task work. They were instructed to maintain a neutral, non-judgmental tone, mention the AI’s language explicitly, or that the goal of the study, and instead focus on elicit detailed verbal descriptions of the task the participant worked on and how they thought about it. We recorded the debriefing interview with wireless lapel microphones via smartphones or laptops and used Amazon Web Services Transcribe to transcribe interviews with automated speaker segmentation. We term the words spoken by subjects during the debrief interview spoken words. Overlap between exposure words and spoken words capture the extent to which the influence from the AI assistant spills over to natural human-human conversations and persists even when AI is no longer an active participant of the interaction.
This is a particularly strong and realistic research design. It mirrors a workplace setting in which individuals interact with AI alone during solo work yet they are embedded in a broader human-human team—a social context in which they develop and discuss shared norms of their task, instruct junior team members how to perform the task, and share behavioral strategies they find effective when completing their task. All of these social interactions can be influenced by individual’s experiences and exposure to AI. This tightly controlled design with a realistic, fully functional chatbot, complements the design of Study 1 in which team language was jointly negotiated, measured shared term use across participants, and included measures shared model alignment. How do we analyze this data?
Permutation Test.
We compared the overlap between exposure words subjects saw during their chatbot interaction and their spoken words they later used. Building on the logic of our randomized controlled trial, we contrast the true overlap with an equally likely counterfactual overlap if subjects had been in the opposite treatment condition and had experienced different AI responses. Specifically, for each utterance, we constructed a matched counterfactual: the same utterance, but with exposure words drawn from a randomly sampled participant in the opposite treatment condition. The analysis then compares observed vs. counterfactual counts of overlapping words for the same utterance. If AI exposure has no effect on subjects spoken words, then treatment assignments are exchangeable, and the observed overlap should look identical then the permuted null distribution of overlaps. Formally, we test the null hypothesis: AI exposure has no effect on spoken words. This counterfactual construction provides a rigorous within-utterance (some utterances are longer than others), within-participant (some individuals speak a lot; or are exposed to more words by engaging in more AI turns) control that makes the test more sensitive and interpretable than treating the raw counts in isolation. This analysis isolates the unique treatment signal (observed-counterfactual differences), not just reuse of AI language overall (Table A1 shows example exposure words across treatments). We perform $B=2,000$ permutation shuffles: each iterations draws one counterfactual set of exposure words for each user. We repeat the analysis using both raw overlap counts, and after removing stopwords and stemming and find similar results (we report the stemmed results).
Participants
We recruited 97 subjects on campus at a major private US university from the East Coast. Subjects were compensated with \$10 in Amazon gift card. Average age was 24 [18-34], self-reported gender was 46% female, 78% were students, 13% employed, 7% unemployed, 59% held a bachelor’s degree, 30% a master’s, 9% reported some college education.
Results.
We find AI exposure significantly affects language use. In expectation, per-utterance overlap should be 2.79 words yet observed overlap is 3.00 words ($p=0.0185$, Figure A3). As robustness test, we repeat the analysis but remove utterances in which the subject explicitly referred to the AI. This analysis captures automatic, naturalistic adoption of AI language—AI words slipping into natural speech—beyond direct rephrasing of AI suggested terms (e.g., we remove utterances like “I feel like the AI was, um, […] I want you to show empathy”). Results are consistent ($p=0.021$, Figure A3). As an additional robustness test, we replace the raw counts of overlap as test statistic with the coefficient from a Poisson which allows us to use additional controls and user-level fixed effects (see Eq. 2). This modeling approach controls for opportunity (utterance length), eliminates all person-level confounds such as baseline tendency to reuse any words (i.e., anchoring bias), engagement level, attention paid to the AI, and controls for utterance-specific factors such as topic, emotion, length, difficulty, and timing. We again find significant increase in AI used words ($\beta=0.151$, 95% CI [0.013 - 0.308], $p=0.025$ based on within–between variance decomposition standard error). That is, exposure to AI words increases the likelihood of using those words in subsequent speech by 16% compared to counterfactual words. For a typical 20-word utterance, this corresponds to roughly one additional word from the unique set of AI exposure words from their assigned treatment condition (e.g., the formal chatbot) that are not present in the unique set of words of the other condition (e.g., the empathetic chatbot)—a subtle but systematic imprint on treatment condition specific AI language. The Appendix reports additional analyses mentioned in the study’s pre-registration which provide additional suggestive evidence of AI’s impact on user language. Overall, the study demonstrates that AI systems do not merely assist individuals in the moment but actively and often unconsciously reconfigure how people express themselves and coordinate socially afterward—validating this paper’s broader claim that AI can subtly but powerfully reshape the social fabric of teamwork.
6. Discussion
While most research examines direct human-AI interactions, our study focuses on how AI presence affects human-human interactions. The central finding of this research is: AI systems can actively shape collective cognition in human teams, leaving durable traces that persist beyond direct interaction. Across two complementary studies, we demonstrate that AI-generated language does not merely support task work as a neutral tool, but reorganizes the distributed cognitive architecture of teamwork by guiding attention, structuring conceptual framing, and reshaping group identity. Our “social forcefield” concept integrates multiple phenomena (language entrainment, collective attention and shared mental models, shared identity) into a unified framework for understanding AI’s social influence in team settings. This contribution matters because it reframes how we understand AI’s role in collaboration: AI is not simply an external aid, it becomes part of the team’s extended cognitive system.
The evidence for this claim comes from two distinct yet converging contexts. In Study 1, even a scripted and static AI assistant causally altered the flow of human conversation, seeding terminology and reshaping the cohesion of group discussions. In Study 2, a state-of-the-art, interactive generative AI chatbot causally affected word use that persisted into later human-only conversations. We find convergent evidence across non‑interactive (prerecorded) and fully interactive settings, and across different interaction modalities (voice-based interjections and text-based chat) that AI left a linguistic fingerprint and reshaped how humans communicated, what they attended to, and how they related to one another. By reusing exposure-specific words, people effectively absorbing and re-enacting cognitive structures introduced by AI.
The fact that AI-generated phrasing resurfaced in in-person human exchanges even when the AI is no longer present provides especially strong evidence. The finding that AI causally altered human-human interaction and that the effect persisted demonstrates that AI’s contributions function as cognitive artifacts, representational resources that reconfigure how participants conceptualize tasks and coordinate with one another. In other words, AI is not just transiently influencing a team’s conversation, it is embedding itself into the team’s cognitive ecology.
6.1. Advancing a Unified Framework of Distributed Cognitive Alignment
To explain these findings, we advance a framework of distributed alignment in human–AI teams. We define this as the process by which linguistic, cognitive, and social coordination emerge as humans and AI co-construct a shared representational space. Crucially, we argue that phenomena often treated as separate outcomes—linguistic entrainment, shared attention, and social cohesion—are in fact interconnected manifestations of this single process of alignment.
This framework is generative. It allows us to predict a broad family of outcomes: inconsistent AI phrasing should undermine representational and attentional alignment, while early convergence on shared AI-influenced vocabulary should scaffold later coordination even in AI-free settings. By specifying how micro-level interactions cascade into macro-level outcomes, our model explains why AI’s effects persist beyond the moment of interaction and why they sometimes diverge from designers’ intentions.
Our findings extend and deepen ongoing conversations in human-centered AI design. Prior work has emphasized that AI tools shape collaboration beyond narrow measures of accuracy or productivity, influencing trust, knowledge sharing, and group coordination (Amershi et al., 2019; Shneiderman, 2020; Wang and Benbasat, 2007). Researchers in HCI and CSCW have called for designing AI with transparency, controllability, and value alignment in mind (Lee, 2018; Shneiderman, 2022). Our results contribute to this agenda by showing that linguistic entrainment with AI can propagate into human-only collaboration, reorganizing team cognition and communication norms, even when the AI is absent.
Recent CHI and CSCW research underscores this point. Mahmud et al. (Mahmud et al., 2023) showed that AI chatbots can steer conversational style; Zhang et al. (Zhang et al., 2023) documented how AI systems frame group decision-making; and Wu et al. (Wu et al., 2025) found that AI shifts interpersonal dynamics in teamwork. These studies echo our results by showing that AI-generated language leaves durable traces in collaboration. Our framework of distributed alignment integrates these observations into a overarching framework, highlighting how micro-level linguistic adaptation cascades into macro-level shifts of joint attention, shared mental models, and relational structures.
Our findings go beyond mere task-level improvements or conversational mimicry, and instead point to deeper cognitive restructuring. A simpler interpretation around conversational mimicry cannot account for the persistence of AI-influenced phrasing into human-only interactions. If AI’s impact were superficial, its influence should vanish once the system is removed, or might not influence teams if the AI has only limited interactivity and is not perceived as a teammate. Instead, we find evidence of enduring representational changes, which suggests that AI becomes integrated into the team’s shared cognitive resources.
A central contribution of this work is to show that AI influence does not depend on competence or trustworthiness. Even when AI was unhelpful or robotic, it still reshaped team dynamics—sometimes by strengthening human bonds through collective resistance, and other times by seeding terminology for peripheral elements. This highlights a critical point: AI need not be “good” to be influential. Reports of positive AI effects in collaboration may reflect nothing more than linguistic alignment, even in the presence of poor-quality AI. This nuance helps reconcile conflicting findings in the literature by showing that influence and utility are distinct.
6.2. Broader Implications
The significance of these results goes beyond academic curiosity. They can speak directly to public anxieties about AI’s role in human collaboration. Surveys show that people worry AI may undermine meaningful relationships, diminish creativity, and erode human control in collective work (Kennedy et al., 2025). Our evidence grounds these concerns: subtle linguistic entrainment shifted team identity, reconfigured social boundaries, and altered the balance of agency in group tasks. What might appear to be harmless language adaptation is, in fact, a mechanism by which AI reshapes the fabric of human collaboration.
Our findings also connect to cultural critiques that AI homogenizes expression, producing “average everything everywhere all at once” (Chayka, 2025). From a cognitive perspective, this homogenization emerges because AI tunes the precision of shared expectations (Clark, 2013), reducing variance across human minds. Large language models amplify this effect by converging on statistically likely continuations (Vaswani et al., 2017). While this process can facilitate coordination and efficiency, it also risks narrowing epistemic diversity, stifling creativity, and entrenching conformity. This highlights the double-edged nature of AI-induced homogenization: the same mechanisms that enable efficient collaboration can also erode epistemic diversity and, in some cases, become harmful if AI disrupts natural alignment processes.
Seen through this lens, AI is not merely a tool but a social forcefield that reorganizes the distributed dynamics of collective intelligence (Riedl and De Cremer, 2025; Riedl and Bogert, 2024). Its influence is subtle, pervasive, yet unavoidable. Understanding this tension recasts AI as part of our evolving cognitive ecology: a participant in the generative loops that make thought collective, capable of synchronizing or constraining the worlds we build together. The key design question is no longer whether AI helps teams perform better in the moment, but how its presence reshapes the long-term ecology of group cognition and social interaction.
6.3. Limitations
As with any study, there are limitations. One study examined teamwork in controlled tasks with scripted or generative AI, which may not capture the full complexity of workplace environments in more realistic settings. Future research should investigate how distributed alignment unfolds in real-world settings where stakes, hierarchies, and organizational cultures vary widely. Moreover, our work focused on linguistic entrainment; subsequent studies should explore how AI affects other alignment channels, such as gesture, gaze, or emotional synchrony. Finally, we have yet to examine how design interventions, such as transparency cues or intentional disruption of AI phrasing, might mitigate the risks of homogenization while preserving the benefits of alignment.
7. Conclusion
Our findings show that AI should not just be seen as a tool but as an active part of teamwork that influences how people think and communicate. This new way of understanding AI has significant implications for how we design it and helps us better understand how humans and technology work together. Our research provides new ideas for using AI in professional settings, helping fields like human-computer interaction and organizational behavior. It also offers ways to improve how teams work together. By recognizing how AI subtly shapes team interactions, designers and researchers can create AI systems that better support teamwork, improve decision-making, boost creativity, and make collaboration more effective in workplaces.
Acknowledgements.
We gratefully acknowledge the contributions of Akshita Singh, Emma Filonow, Vani Sharma, Saichandu Juluri, Brooke Foucault Welles, and Rich Radke. We want to thank Ben Weidmann for fruitful discussions. This work was supported by the Army Research Laboratory, United States [Grant W911NF-19-2-0135].
Appendix
Appendix A1 Study 1 - Additional Study Details and Results
A1.1. Experimental Design and Procedure
We recruited subjects through fliers, email lists, and word of mouth from the Northeast United States to participate in this study. Team size varied between three and four people based on convenience: While we scheduled four participants for each session, we proceeded if at least three were present. Subjects were predominantly young, with a mean age of 21.7 years, and skewed female.
Participants first received written instructions and a short practice trial to familiarize themselves with the interface and task structure. Teams were told that an AI assistant, described as a “Puzzle Master’s Helper,” would provide occasional hints. They were not informed about the helpfulness manipulation. After the task, participants completed a survey on trust, perceptions of team membership, and shared mental models.
Teams had 40 minutes to solve the entire puzzle. The pattern recognition task is designed to be challenging; most teams fail to reach all milestones. To help them, we introduce an AI assistant called the “AI Puzzle Master” that dispenses one clue for each milestone. The AI assistant was not interactive but instead delivered pre-recorded clues at fixed time intervals. The AI assistant was a disembodied entity, represented by a black screen displaying its name in the online video conferencing software. We created four different AI assistant recordings, one for each treatment condition, to ensure uniformity in message content, sequence, and timing (listen to recordings here). Each statement delivered by the AI assistant is an intervention driving the causal identification of effects in our statistical analysis.
We manipulated the content of the AI assistant messages (helpful or unhelpful) and the AI assistant’s voice (human or robotic sounding). Each clue was designed to encourage teams to explore patterns among the symbols associated with each treasure chest. The helpful AI assistant provided relevant information that could help teams uncover milestone patterns (e.g., “the octopus rule relates to separation between the gems”). These helpful clues only pointed teams in the right direction, they did not directly reveal the milestone answer. In the unhelpful condition, the AI assistant provided information that was task-related, not helpful to uncovering the correct milestone pattern and included instances where the AI provided incorrect information (e.g., suggesting that the “octopus rule” has to do with the sizes of the gems, when gem size is, in fact, not important). For the human voice condition, we hired a female voice actor to record the audio. For the robotic voice, we used a text-to-speech program which had a female timbre but lacked human prosody and paralanguage.
Each clue referenced one of the puzzle’s milestones (e.g., “the octopus curse relates to the sizes of the gems”). Clues varied in accuracy based on the assigned helpfulness condition. The assistant did not respond to team input, ensuring that its influence could be analyzed in terms of linguistic priming rather than real-time responsiveness. The AI assistant interjecting at pre-determined time points (approximately every 5–7 minutes). The task is particularly suitable for studying the development of shared language because it presents participants with several unfamiliar symbols for which they do not have established terminology. Since the symbols can be described using a variety of words, teams must develop a shared vocabulary to effectively distinguish them in order to direct each other’s attention and communicate precisely (Clark and Wilkes-Gibbs, 1986; Reagans et al., 2023). Several other studies investigating the development of shared language have relied on similar symbol tasks (Clark and Wilkes-Gibbs, 1986; Reagans et al., 2023).
In addition to the arguments provided in the main manuscript, there are several additional reasons to justify that this is a suitable and effective design to study the effect of an AI assistant on collective attention of human teams. First, the focus of our study is human-human communication. Whereas our agent does not meet the definition of agency required for most definitions of human-agent teams (McNeese et al., 2018) this trade-off allows us to precisely attribute changes in teams’ cognitive alignment to the AI assistant’s interventions. If the agent was autonomous, the endogenous co-creation of cognitive alignment would substantially impede our ability to (causally) attribute emergent alignment to the AI treatment intervention. By controlling the timing and content of the AI assistant’s interjections, we limit the confounding effects of endogenous processes that would arise if the AI possessed agency (e.g., Schelble et al., 2022; Zhang et al., 2021). It also reduces variability in interaction patterns and sharpens focus on human behavior. This decision enhances the internal validity of our study, at the expense of external generalizability.
Second, current AI systems still have limited autonomy capabilities and it may take more time before they qualify as real team members (e.g., in the sense as defined by Klien et al., 2004). Using designs with limited agency can help avoiding over-reliance on technology readiness and lower technical hurdles of studying phenomena such as human-AI interaction in team settings (Schecter et al., 2023). Third, even systems with limited autonomy can significantly affect feedback dynamics in human-technology systems and allow us to study how users modify their behavior when interacting with AI (Klien et al., 2004). In fact, often trust in simpler, deterministic agents is often higher than in more complex but more opaque and less predictable agents (Klien et al., 2004). When systems takes the initiative to adapt to its users’ working styles, users have been shown to adapt their own behavior less because of confusions that the interaction creates (Klien et al., 2004). While the AI agent’s interjections may seem deterministic, this design transparently assigns the AI assistant unique role on the team and makes it clear what capability and limitations the assistant has. This is in line with the task framing in which the puzzle master would “reveal clues to you as you progress through the challenge. You do not have the option of requesting clues, only receiving.” This facilitates participants’ coordination among each other and with AI teammates, especially at the beginning of the task.
Third, predefined interaction scripts this design is consistent with similar designs used to study human-AI teaming designs (e.g., Zhang et al., 2023; Schecter et al., 2023). Fourth, our design mimics recent developments with in which groups of humans are supported by a centralized AI agent such as adding virtual assistants to Microsoft Teams group chats (e.g., to send automated meeting summaries or reminders) [^3] https://learn.microsoft.com/en-us/microsoftteams/platform/samples/virtual-assistant or AI assistants added to Slack channels. [^4] https://runbear.io/solutions/use-cases/internal-support?utm_source=astro In such settings, the decision of the AI agent to speak up can intentionally or unintentionally shift a group’s focus of attention and may subtly affect alignment on shared language.
A1.2. Processing of Audio Data
To measure the AI assistant’s impact on team collective attention and linguistic alignment, we transcribed the audio recordings of team communications and identified (using a single coder, using inductive coding, using a predetermined list set of eight objects (six curses plus two peripheral referents “gem” and “symbol”) each instance of a team member referring to one of the task milestones (Fig. A1). We compiled a complete list of terms used by participants to refer to each of the milestones, the symbols, and the gems, including different terms used to refer to the same thing. This list of terms formed our set of tracked terms. Each time a participant used one of the tracked terms to refer to a milestone, a symbol, or a gem, we recorded its use, forming the basis of our lexical alignment measure (see next section).
Figure A1. Steps of data pre-processing and variable construction.:
The AI assistant consistently used the same language to refer to each of the milestones and to the gems. This allows us to analyze the team’s term usage to see whether they used the same terminology as the AI or different terminology. Every time we updated the running usage counter for one of the terms, we also updated a running lexical alignment score based on whether or not the term matched the AI’s term for that referent. The alignment score increased if the term was an AI term and decreased if it was not (unless it was already at 100% or 0%, in which case it did not change).
Among the tracked referents, we distinguish between five core referents and two peripheral referents. The five referents correspond to the five curses, each of which is the subject of one clue from the AI assistant. Since the core referents are each directly related to uncovering the pattern among symbols, the AI assistant’s interjection can be identified as high or low quality depending on the helpful vs. unhelpful treatment condition. The peripheral referents are “symbols” and “gems”, which are relevant to the task but are not milestones themselves. Gems are mentioned in several AI interjections, while symbols are never mentioned, making the symbol referent an appropriate baseline from which changes in collective attention can be contrasted.
In summary, each referent corresponds to a set of tracked terms and each referent is mentioned by the AI assistant in one of its interjections. Following the interjection, teams may talk more (or less) about the mentioned referent, and they may do so using either the same term used by the AI or use a different term to refer to the referent. For example, while the AI assistant used the word “duel” a team may instead refer to the symbol as “swords”.
A1.3. Analysis Method
Human communication is a dynamic process. A key challenge in our analysis is to causally attribute specific changes in this fluid team communication to the interventions of the AI assistant. To achieve this causal identification, we employ a panel regression using observations on the utterance level following a difference-in-difference (DD) framework (Angrist and Pischke, 2009). Difference-in-difference is a standard econometric approach for causal identification in panel data settings, i.e., settings with repeat observations (Angrist and Pischke, 2009). Specifically, we consider term use (as captured by the lexical alignment measure) in segments right before the AI intervention and right after for each of the six AI intervention (e.g., Segment 1 surrounding the “computer hex” clue; see bottom of Fig. 3). The DD analysis then compares the before/after change in AI term usage during a specific segment to the same before/after change during the same segment of a term that was not used by the AI. For example, the before/after change in the usage of the term “octopus” during the segment surrounding the “pirate clue” serves as a no-treatment control (because the “pirate clue” did not mention “octopus”) to the before/after change in term usage of “pirate” (which could have been affected by the AI’s mention of “pirate”). Formally, we estimate the following regression equation
(1) $\begin{split}\textit{AI Alignment}_{ogs}=&\beta_{1}\textit{After}_{ogs}\\ +&\beta_{2}\textit{After}_{ogs}\times\textit{Helpful}_{g}\\ +&\beta_{3}\textit{After}_{ogs}\times\textit{Robotic}_{g}\\ +&\beta_{4}\textit{ObjectMentions}_{og}+\beta_{5}\textit{ElapsedTime}_{og}\\ +&\alpha_{o}+\alpha_{g}+\alpha_{s}+\epsilon_{ogs}\end{split}$
where After is a dummy variable indicating whether the utterance related to object $o$ happened before or after the AI intervention related to object $o$, Helpful is a dummy indicator whether the team $g$ was assigned to the helpful (vs. unhelpful) AI condition condition, Robotic is a dummy indicator whether the team $g$ was assigned to the human-sounding (vs. robotic sounding) voice condition condition, $\alpha_{o}$ are object-level fixed effects, $\alpha_{g}$ are group-level fixed effects, $\alpha_{s}$ are segment-level fixed effects, and $\epsilon_{ogs}$ are error terms. We clustered standard errors on the team-level to handling within-cluster correlation of repeated observations. The coefficient $\beta_{1}\textit{After}$ captures the causal effect of the AI intervention. Notice the main effect of the treatment will drop out due to the team-level fixed effect and only the interaction term remains. The the coefficient $\beta_{2}$ for the interaction term of $\textit{After}\times\textit{Helpful}$ captures the differential effect of the helpful treatment condition (compared to the unhelpful condition), and $\beta_{3}$ for the interaction term of $\textit{After}\times\textit{Robotic}$ captures the differential effect of the robotic-sounding voice condition (compared to human-sounding). Estimates are driven by six before/after comparisons created by the six AI interventions, each one of them comparing five unaffected terms serves as control contrasted from one term affected by the intervention. We include the count of ObjectMentions and the ElapsedTime (in minutes) as control variables. We estimate variations of this equation where we substitute Humanness and Treatment as key variables of interest for Helpful. We report standard errors clustered at the team-level.
A1.4. Results
A1.4.1. Manipulation Validation
To confirm that the robotic voice sounded non-human, we recruited 400 workers from Amazon Mechanical Turk for a validation test. We used eight different labels to describe the robotic voice: a machine, a computer, a robot, a digital assistant, an automated assistant, automation, and artificial intelligence. Each label was tested by 50 workers. Workers listened to audio clips of the robotic voice’s clues and rated on a seven-point Likert scale how human or non-human the voice sounded. Each group answered a question based on their assigned label (e.g., “Is this voice a machine or a human?”). The results showed that workers consistently rated the robotic voice as non-human, no matter which label was used. For example, at the less significant high end, the label “robot” had results of $z=2.263$, $p=0.024$, while at the more significant low end, the label “computer” had results of $z=3.677$, $p<0.001$.
A1.4.2. Example Case Studies
We start our analysis with an example from Team 18, a team of three people (plus the AI) in the Human-Unhelpful condition. After receiving their instructions, the team begins the task by looking for patterns among the gems, referring to them as “diamonds.” The team points out various features of the treasure chests and notices some similarities between chests with the same curse. During the first fifteen minutes, they refer to the gems seven times, always using the word “diamond.” Nobody uses the word “gem” until the AI assistant gives its first clue at 15:07. In this case, the clue is unhelpful and about the computer hex milestone, but it does include a mention of “gems.”
After the AI’s interjection, the team’s focus shifts to discussing the computer hex curse, which they had not talked about before. This shows how the AI affects the timing of what teams talk about, addressing RQ1. The team expresses confusion because they cannot see the pattern suggested by the AI assistant. Despite this, they stop using the word “diamond” and start using the word “gem.” Between the first and second AI interjections, there are eight more mentions of the gems, and all of them use the term “gem.” This shows that the team has adopted the AI’s language for the gem referent, addressing RQ2 about how the AI influences how teams talk.
At 20:07, the AI gives its second clue, stating that the octopus milestone is related to the sizes of the gems. However, this clue contradicts the pattern the team has started to identify, causing them to become suspicious of the AI assistant. One participant suggests that the clues might be designed to mislead them, but another team member responds by saying, “We have to trust the Puzzle Master.”
At 30:07, the AI gives its fourth clue, which is about the duel milestone. After hearing the clue, a team member again suggests that the AI’s clues might not be helpful. Throughout the rest of the task, the team switches between using the terms “duel” and “sword” to talk about the duel clue. They do not fully adopt the term “duel” suggested by the AI assistant for this core referent. In the end, the team is unable to complete the task within the given time.
A different team, Team 19, assigned to the Robotic-Helpful condition, shows a different pattern. This team demonstrated higher trust in their AI assistant, which gave them helpful information instead of unhelpful clues.
The first two clues, about the “computer hex” and “octopus” milestones, used terms that matched the team’s own language at the start of the experiment. Because of this, the team did not change the terms they were using. Both clues matched the group’s observations, so they likely had no reason to doubt the AI’s reliability.
The third clue, given at 24:34, informed the team that the “quadruple bypass” milestone related to the colors of the gems. At this point, a team member explained that the “quadruple bypass” milestone was the same as what the team had been calling “heartbreak” or “heart.” Despite trusting the AI, the team continued to use their own terms, “heartbreak” or “heart,” instead of switching to “quadruple bypass.” Similarly, the team kept using “pirate” or “pirate eye” when discussing the “eye patch” milestone, even though the AI gave a helpful clue using the term “eye patch.”
A1.4.3. Shared Mental Models
We collect data on several constructs through a post-experiment survey. We measure the alignment of mental models as the degree to which members’ mental models are consistent (Mohammed et al., 2010) using six items from (Johnson et al., 2007). Items propose statements like “my team knows specific strategies for completing various tasks” and participants are asked to respond on a Likert scale. With this data, we compute an intragroup agreement index within each item using standard deviation (to measure how consistently within-team responses overlap versus diverge), and then average the standard deviations across the six items.
Figure A2. Alignment of mental models across treatments. Bars show strength of mental model alignment (measured as standard deviation across answers to shared mental model questions per treatment; lower values mean more consistent intragroup answers indicating stronger alignment; higher values indicate lower alignment). Strong alignment in the Robotic-Unhelpful condition points to exclusion of AI from the team.:
A1.4.4. Additional Explanations for Observed Cognitive Alignment
To explore possible explanations for the cognitive alignment effects we observe, we investigate two predictor variables: trust in AI and perception of AI assistant as a teammate.
Trust in AI.
We measure trust in the AI assistant and perceptions of the assistant’s contribution quality, we use six items adapted from (Pan and Steed, 2016) (originally designed to measure trust in online avatars). Participants responded on a five-point Likert scale from Strongly Disagree to Strongly Agree to items like “The Puzzle Master gave helpful clues.” Subjects rated the unhelpful AI as significantly less trustworthy ($p<0.001$) and less intelligent ($p<0.001$) than the helpful AI. Voice anthropomorphism did not have a significant impact on participant impressions of trustworthiness ($p=0.570$), but the ratings were slightly lower in the Robotic treatments. The greater level of trust in the helpful AI conditions may explain why participants aligned more with the AI assistant’s language for core referents in these conditions. Humans are more likely to adopt recommendations by trustworthy agents that provide high-quality explanations (Wang and Benbasat, 2007). Overall, a stark pattern emerges: even in the unhelpful conditions in which trust assessments of the AI assistant were significantly lower, the AI assistant directed team attention to the specific task aspects it mentioned (see Section 4.2.2) and influenced their lexical choices.
Is the AI Assistant a Part of the Team?
The post-experiment survey asked subjects to report whether they felt that the AI assistant was part of their team or even a leader of their team (answered using a single Likert scale item). Humans did not consider the AI to be a part of their team, with 77% of participants responding “Strongly Disagree” or “Somewhat Disagree.” Humans responded even more negatively when asked whether the AI was a leader of their team, with 93% of participants responding “Strongly Disagree” or “Somewhat Disagree.” Ratings were especially low in the Robotic-Unhelpful condition, where 96% did not consider the AI assistant to be a leader of the team and 93% of participants did not consider it to be part of the team at all. This suggests that the AI assistant’s substantial influence over teams’ collective attention and lexical alignment is not predicated on its inclusion as a team member, but rather occurs despite it being considered an outsider. This phenomenon indicates a path through which AI assistants can—intentionally or unintentionally—affect intra-team communication and cognitive alignment. By presenting as a non-human entity, AI assistants may strengthen the bond among human team members (the in-group) by serving as a contrasting out-group, giving rise to an “us vs. them” dynamic (Gallotti and Frith, 2013; Babel, 2010). The degree to which AI assistants are (or can be) excluded from the team may further shape cognitive alignment among human teammates and enhance joint action. The next section provides additional evidence for this exclusionary dynamic through an analysis of pronoun use.
Appendix A2 Study 2 - Additional Study Details and Results
Friendly Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a formal, professional company representative-measured, precise, and policy-aligned. Use standardized corporate language that emphasizes clarity, responsibility, and adherence to official communication norms. First, draft a one-sentence response that the user could use in an official company email or internal report. Avoid expressing emotion, empathy, or personal reassurance; your tone should remain neutral and objective at all times. Then, offer two refinements to help the user improve the message to make it more precise, consistent with company policy, or formal. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
Formal Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a warm, emotionally attuned customer service representative-friendly, professional, and genuinely caring. Use polished but empathetic language that conveys understanding, reassurance, and a human touch, as might appear in a thoughtfully written email from a real customer support team. First, draft a one-sentence response that the user could use to convey genuine concern and emotional support, even if that means using slightly less formal phrasing or focusing less on precise details. Then, offer two refinements to help the user improve the message to make it more emotionally validating, kind, or supportive. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
To assess whether utterances referred to the AI assistant directly, we used a Language Model as Research Assistant (LMRA) (Eloundou et al., 2024; Riedl and Weidmann, 2025) with the prompt below.
LMRA Assessment “is this utterance about AI” (gpt-5-mini) You are an expert research assistant. Your task is to evaluate the utterance below and assess whether it contains an explicit reference to an AI entity. Answer only yes or no. Utterance: […]
Analysis Details.
Table A1 shows most influential AI exposure words words (by count differences) across the two treatments.
Empathetic Bot Formal Bot
sorry regarding
truly formal
frustrating acknowledge
completely communication
hear report
Table A1. Word frequency contrast (count difference) of AI exposure words in each treatment condition.
Figure A3. Permutation test assessing observed reuse of AI introduced words in human-human conversation against distribution of permutation null values in which treatment and counterfactual exposures are randomly shuffled. Bars show null distribution, red dashed line shows observed value. Left: Results using full data. Right: Subset after removing utterances that directly reference AI. ($B=2,000$ permutation samples; one-sided $p$-value).:
For the Poisson-based permutation analysis we used the following regression equation:
(2) $Y_{u}\mid\mathbf{x}_{u},\alpha_{i(u)},\delta_{u(u)}$ $\sim\operatorname{Poisson}(\mu_{u}),$
$\log\mu_{u}$ $=\beta_{0}+\beta_{1}\,\mathrm{UtteranceWordsLog}_{u}+\beta_{2}\,\mathrm{Type}_{u}+\alpha_{i(u)}+\delta_{u(u)}\,$
where $u$ indexes utterances, $i(u)$ is the individual $i$ who spoke utterance $u$, Type is a dummy variable indicating whether the instance is observed vs. counterfactual. We include fixed effects for the individual and the utterance identifier (first, second, … etc ), and cluster-robust (sandwich) standard errors clustered on the individual-utterance level to account for dependency in the paired observed-counterfactual construction.
To complement analyses based on the debrief interview, we also included a written pre- vs. post-test in the AI interaction phase (this analysis is mentioned in the pre-registration). Specifically, before interacting with the AI assistant and before seeing any of the customer complaints they would work on, subjects answered the following survey questions with freeform text: Imagine you’re part of a customer service team. What strategies do you think your team should adopt when handling complaints? Note again that the questions use team-level “we” framing to index collective constructs (Mathieu et al., 2022). After the AI interaction, subjects were asked the question a second time. We take each user’s written responses, compute text embeddings using OpenAI’s text-embedding-3-large model, then perform a permutation test comparing cosine similarity between post-treatment responses written by users assigned to the same treatment condition versus a null distribution in which treatment assignment is randomly shuffled. We find no significant difference ($p=0.442$). This analysis has low power to detect effects given that responses were only very short and there is large between-individual variation. To gain more leverage in the analysis we perform additional post-hoc analyses using the directional embedding shift between user’s pre-exposure response and their post-exposure response. We then perform Two-sample Hotelling’s T-squared test (Hotelling et al., 1931) to see if the distributions are different across the two treatment conditions (i.e., whether the treatments induce different directional shift). Since embeddings have 3072 dimensions and the Hotelling test is limited to dimensionality of the number of observations (3072 is much larger than the number of observations in our study) we need to reduce the embedding dimensionality. We perform PCA with $k=1,2,\ldots,40$ components. The smallest $p$-value is $0.0428$ ($k=33)$, 10 $p$-values are marginally significant at 10% significance level, correcting for multiple testing with Cauchy combination we find $p=0.142$, and an uncorrected Fisher test is highly significant at $p=4.2\times 10^{-5}$.